OpenStudy (anonymous):

An identification code is to consist of three letters followed by four digits. How many different codes are possible if repetition is permitted?

OpenStudy (anonymous):

\[26\times 26\times 26\times 10000\]

OpenStudy (anonymous):

i get 175760000

OpenStudy (anonymous):

ok im confused now

OpenStudy (amistre64):

7 spaces to fill: - - - - - - - 26 letters can fill the 1st space; 26 outcomes 26 letters can fill the 2nd space; 26 * 26 outcomes 26 letters can fill the 3rd space; 26*26*26 outcomes 10 numbers can fill teh 4th space; 26*26*26*10 outcomes 10 numbers can fill teh 4th space; 26*26*26*10*10 outcomes 10 numbers can fill teh 4th space; 26*26*26*10*10*10 outcomes 10 numbers can fill teh 4th space; 26*26*26*10*10*10*10 outcomes total possible outcomes? 26^3 * 10^4

OpenStudy (amistre64):

yeah, i copy pasted my typo lol